# A practical guide to Digital Micro-mirror Devices (DMDs) for wavefront shaping

### v2

@@ 2.1 A 1D model
lies in the geometry of the pixel surface.
@@
both the modulation quality and system efficiency.
The impact of these diffraction effects is highly contingent on several factors:
the wavelength of the illumination, the pixel pitch, and both the incident and outgoing angles.
Therefore, alongside selecting an appropriate anti-reflection coating,
@@
with differing intensities
@@
where is the wavelength of the light,
is the angle of the first-order diffraction,
@@
In essence, the incident angle serves to shift the entire angular diffraction pattern
@@
—the tilt angle of the mirror in the DMD projected onto the axis we consider—
@@
and the intensity nears zero for all other orders.
@@
a shift in the envelope and in the position of its maximum,
now indicated by .
@@
We represent in Fig. 3 the angular response of a flat grating and a blazed grating for a 1D filling fraction of 95% (correponding to a 2D filling fraction of %). For a flat grating, the zero-th order contains most of the intensity, the other orders being negligible in comparison. For the blazed grating example shown, we are in a situation close to the worst case scenario: Two diffraction orders have a significant and comparable intensity, and other orders also have a non-negligible contributions. In the optimal scenario, where the peak of the envelope corresponds to a diffraction order, it results in a single diffraction order carrying the majority of the energy. This state is achieved when the conditions of the blazed grating equation are fulfilled [19]:
We note that the incident angle olso affects the diffraction efficiency.
@@ 2.2 The 2D case
Pixels are uniformly repeated along these axes.
@@
A straightforward and prevalent solution is to rotate the chip by 45 degrees
@@
which aligns the pixel’s axis of rotation to be vertical.
@@
The 2D array can be seen as 2 orthogonal 1D gratings in the and directions, both having the same properties. A more comprehensive depiction of the 2D system can be found in [14]. In order to utilize Eq.(1), one first needs to project the different angles of the problem onto the incident planes of the two 1D gratings. This yields , , and , where represents the true rotation angle of the mirrors relative to the diagonal of the pixels, and and denote the incident and outgoing angles in the horizontal plane. We can quantify how close we are to the ideal case, i.e., when satisfying the blazed equation outlined by Eq.(1), by defining a blazed number as introduced in [20]:
with representing the floor function,
and the modulo 2 operation.
@@
To demonstrate the effect, we conduct a simulation of a DMD using Python (refer to tutorial and code in [20]) with two pixel pitches of µm and µm, under a coherent excitation at nm. Fig. 5 shows the estimated blazed number as a function of the angle of incidence in the horizontal plane, along with the far field diffraction pattern for two distinct incident angles. Fig. 6 shows the blazed number as a function of both the incident angle and the wavelength for pixel pitches of µm and µm. It should be noted that the efficiency of diffraction can be altered by adjusting the angle of incidence. However, its impact is relatively confined within an acceptable angular range that aligns with experimental limitations (i.e. for angles far from ). Far-field patterns are centered around the maximum of the envelope (marked by a yellow cross). We see that for small values of , the pixel pitch of µm leads to a blazed number close . It corresponds in the far field to having one bright order of diffraction close to the maximum of the envelope.
@@ 5 Conclusion
Appendix A: 1D calculation of the DMD diffraction effect
Assuming the effect of the device’s finite size and illumination to be negligible, and situating ourselves within the context of the small-angle approximation, we can represent the field reflected from the device under the influence of plane wave illumination across two systems as follows:
with the rectangular function, representing the finite size of the pixel, defined as:
The intensity as a function of the angle in the far-field is given, up to a homotetic transformation, by the absolute value squared of
The Fourier transform of Eq. 8 can be written as:
We observe that the envelope (right hand term) is maximal for while the effect of the periodicity (left hand term) is maximal for , representing the orders of diffraction.
S.M.P, M.W.M, and R.G.C and acknowledge the French Agence Nationale pour la Recherche grants No. ANR-23-CE42-0010-01 MUFFIN and No. ANR-20-CE24-0016 MUPHTA and the Labex WIFI grant No. ANR-10-LABX-24, ANR-10-IDEX-0001-02 PSL*.

### v3

@@ 2.1 A 1D model
lies in the geometry of the pixel surface:
The pixels of the DMD are not flat, but are mirrors that can be tilted
in two different positions around a rotation axis
along the diagonal of the square shape of the pixels.
@@
both modulation quality and system efficiency.
The impact of these diffraction effects is highly dependent on several factors:
the wavelength of illumination, the pixel pitch, and both the incident and outgoing angles.
Therefore, in conjunction with selecting an appropriate anti-reflection coating,
@@
with different intensities
@@
where is the wavelength of the light, is the incident angle,
is the angle of the first diffraction order,
@@
The incident angle has the effect of shifting the entire angular diffraction pattern
@@
—the tilt angle of the micro-mirrors of the DMD—
@@
and the intensity is comparatively small for all other orders.
@@
a shift in the envelope and in the position of its maximum
@@
We represent in Fig. 3 the angular response of a flat grating and a blazed grating for a 1D filling fraction of 95% (correponding to a 2D filling fraction of %). For a flat grating, the zero-th order contains most of the intensity, the other orders being negligible in comparison. For the blazed grating example shown, we are in a situation close to the worst case scenario: Two diffraction orders have a significant and comparable intensity, and other orders also have non-negligible contributions. In the optimal scenario, where the peak of the envelope corresponds to a diffraction order, it results in a single diffraction order carrying the majority of the energy. This state is achieved when the conditions of the blazed grating equation are fulfilled [19]:
@@ 2.2 The 2D case
The pixels are uniformly repeated along these axes.
@@
A straightforward and common solution is to rotate the chip by 45 degrees
@@
which aligns the pixel axis of rotation to be vertical.
@@
A more detailed description of the 2D system and computation of its response can be found in [14]. However, we can find a simple condition for having most of the energy in one diffraction order by looking at the effect of the tilts of the mirrors when the axis is rotated by 45 degrees compared to the axis of the pixel array (axes and ). We consider a plane wave with an incoming angle contained in the horizontal plane (i.e. orthogonal to the rotation axis of the micro-mirrors). The effect of this angle is represented by a phase term of the form:
with the incoming wavevector and the position in the modulator plane, represented by the coordinaxte .
The grating condition is achieved when the output contributions are in phase for the output angle corresponding to the specular reflection on the tilted mirrors. It corresponds to the plot on the left in Fig.3 for the 1D case. The phase contribution resulting from an output angle , corresponding to the specular reflection on each micro-mirror, reads:
The total phase then reads:
We see that and have the same effect. The grating condition is then achieved when two consecutive pixels, separated by a distance in both the and directions, are in phase. I.e. for
The new blazed grating equation then reads :
with an integer.
Note that this is what one would obtain using Eq. 6 for
blazed grating with a pitch , or a 1D grating rotated by 45 degrees.
Those systems are not equivalent in the general case,
but the grating condition in the 2D case is
similar when considering an incident plane wave in the horizontal plane.
We can quantify how close we are to the ideal case, i.e., when satisfying the blazed this equation, by defining a blazed number as introduced in [20]:
where represents the modulo 2 operation.
@@
To demonstrate the effect, we first conduct a simulation of a DMD using Python (refer to tutorial and code in [20]) with two pixel pitches of µm and µm, under a coherent excitation at nm. Fig. 5 shows the estimated blazed number as a function of the angle of incidence in the horizontal plane, along with the far field diffraction pattern for two distinct incident angles. Fig. 6 shows the blazed number as a function of both the incident angle and the wavelength for pixel pitches of µm and µm. It should be noted that the efficiency of diffraction can be altered by adjusting the angle of incidence . However, its impact is relatively confined within an acceptable angular range that aligns with experimental limitations (i.e. for angles far from ). Far-field patterns are centered around the maximum of the envelope for an output angle (marked by a yellow cross). We see that for small positive values of , the pixel pitch of µm leads to a blazed number close to . It corresponds in the far field to having one bright order of diffraction close to the maximum of the envelope.
To illustrate the link between the blaze number and the efficiency of the diffraction setup,
we compute the ratio of the energy in the brightest diffraction spot to the total diffracted energy
for different values of the wavelength,
with fixed values for the pixel pitch (7.6 µm) and incident angle (normal incidence).
We simulate the DMD with no gap between the pixels (i.e., a filling fraction of 100%)
and with 10 pixels in each direction,
where all the mirrors of the pixels are in the same state.
The results are shown in Fig. 7.
We observe that the maxima (resp. minima) of the diffraction efficiency
correspond to the maxima (resp. minima) of the blaze number.
Note that the actual values of the minimum and maximum diffraction efficiency depend on the parameters of the DMD, such as pixel pitch, filling fraction, resolution, etc.
2.3 Experimental measure of the optimal incident angles
We experimentally characterized a DLP9500 DMD with a pixel pitch of µm at nm. The illumination geometry and DMD orientation are as presented in Fig.4. The experimental setup, shown in Fig.8, consists of an objective (L), a lens (L’), and a pinhole to produce a clean, collimated laser beam. A diaphragm is used to control the size of the beam incident on the DMD, and the reflected light is observed on a screen.
We then illuminate with a collimated monochromatic beam the DMD in its full-on (resp. full-off) state, corresponding to all pixels being rotated by an angle (resp. ), with degrees. The light distribution across the diffraction orders can be assessed directly with the naked eye by placing a screen after reflection from the DMD. The DMD is mounted on a rotating plate with its axis of rotation orthogonal to the table, allowing variation of the angle , as defined in Fig. 4. The reference angle () is determined by auto-collimating the incident beam with the DMD in its idle state ().
The DMD is then rotated until the reflected light is almost entirely on a single diffraction order, with adjacent orders extinguished. The corresponding values of are shown in Fig.9 as red lines, illustrating a good correlation between the computed evolution of the blazed number (Eq.7) and the diffraction efficiency.
@@ 5 Conclusion
Acknowledgments
S.M.P, M.W.M, and R.G.C acknowledge the French Agence Nationale pour la Recherche grants No. ANR-23-CE42-0010-01 MUFFIN and No. ANR-20-CE24-0016 MUPHTA and the Labex WIFI grant No. ANR-10-LABX-24, ANR-10-IDEX-0001-02 PSL*.
L.M, J.C, M.G, and R.G acknowledge Cathie Ventalon for sharing a V-module and for her valuable discussions. L.M, J.C, M.G, and R.G were funded by Centre National de la Recherche Scientifique MITI/80Prime and the iXcore - iXlife - iXblue Fundation.